Problem 2 The Wood Company has the following costs: Category Cost Cost Rent Utilities Insurance Payment Period Monthly Monthlplease answer number two! Thank you
Answer #1

Fixed cost per month = 1500+500+(900/4)+(3600/12) = 2525

a) Break even unit = Fixed cost/Contribution margin per unit = 2525/15-4.75 = 246 Units

Break even sales = 246*15 = 3690

b) Monthly fixed cost 300 Units = $2525

c) Variable cost per unit = 4.75